# Which Similarity-Sensitive Entropy (S-entropy)?

### v3

@@ Abstract
Shannon entropy is not the only entropy that is relevant to machine‑learning datasets, nor possibly even the most important one. Traditional entropies such as Shannon entropy capture information represented by elements’ frequencies but not the richer information encoded by their similarities and differences. Capturing the latter requires similarity‑sensitive entropy (S‑entropy). S‑entropy can be measured using either the recently developed Leinster–Cobbold–Reeve framework (LCR) or the newer Vendi score (VS). This raises the practical question of which one to use: LCR or VS. Here we address this question conceptually, analytically, and experimentally, using 53 large and well-known imaging and tabular datasets. We find that LCR and VS values can differ by orders of magnitude and are complementary, except in limiting cases. We show that both LCR and VS results depend on how similarities are scaled, and introduce the notion of “half‑distance” to parameterize this dependence. We prove that VS provides an upper bound on LCR for several values of the Rényi–Hill order parameter and present evidence that this bound holds for all values. We conclude that VS is preferable only when a dataset’s elements can be usefully interpreted as linear combinations of a more fundamental set of “ur‑elements” or when the system that the dataset describes has a quantum‑mechanical character. In the broader case where one simply wishes to capture the rich information encoded by elements’ similarities and differences as well as their frequencies, LCR is favored; nevertheless, for certain half‑distances the two methods can complement each other.
1 Introduction
@@
Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across not only science and engineering but in commerce and law [15, 29, 22, 31]. Traditional entropy, as formulated by Shannon [30] and generalized by Rényi [28] (and later Tsallis [32]), depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits (or nats) required to encode the shape of the distribution, with skew distributions requiring fewer bits and flatter distributions requiring more (Fig. 1). These entropies rely solely on frequencies and therefore ignore a rich additional source of information about the system: the pairwise similarities and differences among its elements (Fig. 1).
@@
Figure 1: The concepts of element, frequency distribution, traditional entropy (at $q=1$), similarity function, similarity measure ($z_{i}j$, with examples), similarity matrix ($Z$), and S-entropy (here, LCR, also at $q=1$). Entropy values are expressed in effective-number form, i.e. in units of effective number of images present in the dataset. Different similarity measures can be chosen (see Section 5.2); in this example, the similarity measure is the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
@@
Incorporating similarity into the calculation results in so-called similarity-sensitive entropy or S-entropy. This was first accomplished by Leinster and Cobbold [14] and extended by Reeve and colleagues [27], in a framework we refer to as LCR [17]. By incorporating similarity, LCR can differentiate between systems that have identical frequency distributions but whose elements vary in how alike they are (Fig. 1). This capability is valuable in many domains [23], especially in machine learning (ML) [6, 7] (for example, to achieve state-of-the-art performance more efficiently [5]), where datasets generally consist of all-unique elements, meaning any two datasets that are the same size will have the same traditional (i.e. similarity-insensitive) entropies because their frequency distributions are both flat, even though they may differ dramatically in composition. S-entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [1, 2], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat, such that traditional entropies are less informative.
@@
More recently, an alternative form of S-entropy to LCR has been described, consisting of the Vendi score [9, 26] and its variants or “cousins” [25]—hereafter collectively VS. This alternative raises the practical question of which form of S-entropy to choose for a given system or dataset: LCR or VS. The present work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets commonly used for ML research, and comparing their values.
2 Related Work
@@
Entropies fall into two classes: (i) traditional, i.e. similarity‑insensitive, and (ii) similarity‑sensitive, a.k.a. S-entropy. Although the literature often refers to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), both classes are actually families of measures; individual members are distinguished by how they weight frequencies and, for S-entropy, by their similarity matrix as well (see Table 1).
2.1 Traditional, similarity-insensitive entropy
@@
Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the dataset or system [13]. Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ give greater weight to more frequent elements. $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form. Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [12]. For this reason, Hill’s formulation is often described as a unifying framework. $D_{q}$ are known as the Hill numbers or D numbers (for “diversity”) and are given by
$D_{q}(\mathbf{p})=\exp(H_{\alpha}(\mathbf{p}))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}^{q}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln p_{i}\right),&q=1.\end{cases}$ (1)
$\mathbf{p}$ denotes the frequency distribution of unique elements. Other generalizations such as Tsallis entropy [32] also exist. What these traditional entropies have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
2.2 S-entropy
2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
LCR [17] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropies do not. A similarity matrix $Z$ is introduced, with entries $z_{ij}\in[0,1]$ that quantify the similarity between elements $i$ and $j$. The similarity of each element to itself is set to 1, making $Z$’s diagonal entries 1. The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity‑sensitive versions of the Rényi entropies $H_{\alpha}^{Z}$ and are given by
$D_{q}^{Z}(\mathbf{p};Z)=\exp(H^{Z}_{\alpha}(\mathbf{p};Z))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}(Z\mathbf{p})_{i}^{\,q-1}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln(Z\mathbf{p})_{i}\right),&q=1.\end{cases}$ (2)
Here $(Z\mathbf{p})_{i}=\sum_{j}z_{ij}p_{j}$ is the frequency‑weighted average similarity of element $i$ to all the elements (including itself). When this average is large, element $i$ is considered ordinary, making Eq. 2 interpretable as the average “ordinariness” [14] across all elements. $D_{q}^{Z}{(\mathbf{p})}$ appears widely in the recent literature, where it is variously known as:
@@ 2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
  • similarity-sensitive or similarity-aware diversity more generally [23], for which “S-entropy” is a convenient shorthand.
LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become less informative (Fig. 1). Representative applications in the life sciences include high‑throughput immunology (immunomes) [2], microbiome research (metagenomics) [23], and medical imaging [6] Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g. images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [6].
2.2.2 The Vendi score and its variants (VS)
VS constitutes a related but separate class of similarity‑sensitive entropy measures. Like LCR, VS entropies are functions of a similarity matrix, but now the matrix has dimensions $n\times n$, where $n$ is the number of observations or total elements; for this reason we refer to it as $Z_{n}$ to distinguish it from the $Z$ used in LCR, which has one row/column per unique element. (When all elements are unique, $Z$ and $Z_{n}$ are the same.) We can define $Z_{n}$ in terms of $Z$ as
@@ 2.2.2 The Vendi score and its variants (VS)
The original VS is defined as the exponential of the Shannon entropy of the eigenvalues $\lambda_{i}$ of $Z_{n}/n$. The division by $n$ normalizes $Z_{n}$ to unit trace (because as in $Z$, the diagonals of $Z_{n}$ equal 1):
$\mathrm{VS}{(\mathbf{p}_{n};Z_{n})}=\exp\!\left(-\sum_{i=1}^{n}\lambda_{i}\log\lambda_{i}\right)$ (4)
VS is to this similarity matrix what the von Neumann entropy is to the quantum density matrix. Replacing the Shannon term in Eq. 4 with a Rényi entropy of order $q$ yields the so‑called “cousins” [25] of the Vendi score:
$\mathrm{VS}_{q}{(\mathbf{p}_{n};Z_{n})}=\left(\sum_{i=1}^{n}\lambda_{i}^{q}\right)^{\frac{1}{1-q}}$ (5)
@@
$H_{\alpha}(mathbf{p})=\frac{1}{1-\alpha}\log\sum_{i}p_{i}^{\alpha}$
@@
$q=1$ Yes $\lambda_{i}$ are eigenvalues of $Z_{n}=Z/N$ (positive semi-definite). ML diversity, Generative model evaluation, Ecology
@@
$\text{VS}_{q}(\mathbf{x},\mathbf{k})=\left(\sum_{i}\lambda_{i}^{q}\right)^{\tfrac{1}{1-q}}$
@@
3 Methods
3.1 Datasets and definitions

### v4

# Which Similarity-Sensitive Entropy (Sentropy)?

@@ Abstract
Shannon entropy is not the only entropy that is relevant to machine-learning datasets, nor possibly even the most important one. Traditional entropies such as Shannon entropy capture information represented by elements’ frequencies but not the richer information encoded by their similarities and differences. Capturing the latter requires similarity-sensitive entropy (“sentropy”). Sentropy can be measured using either the recently developed Leinster-Cobbold-Reeve framework (LCR) or the newer Vendi score (VS). This raises the practical question of which one to use: LCR or VS. Here we address this question theoretically and numerically, using 53 large and well-known imaging and tabular datasets. We find that LCR and VS values can differ by orders of magnitude and are complementary, except in limiting cases. We show that both LCR and VS results depend on how similarities are scaled, and introduce the notion of “half-distance” to parameterize this dependence. We prove the VS provides an upper bound on LCR for all non-negative values of the Rényi-Hill order parameter, as well as for negative values in the special case that the similarity matrix is full rank. We conclude that VS is preferable only when a dataset’s elements can be usefully interpreted as linear combinations of a more fundamental set of “ur-elements” or when the system that the dataset describes has a quantum-mechanical character. In the broader case where one simply wishes to capture the rich information encoded by elements’ similarities and differences as well as their frequencies, we propose that LCR should be favored; nevertheless, for certain half-distances the two methods can complement each other.
I Introduction
@@
Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across not only science and engineering but in commerce and law [25, 44, 36, 46]. Traditional entropy, as formulated by Shannon [45] and generalized by Rényi [43] and later Tsallis [47], depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits (or nats) required to encode the shape of the distribution, with skew distributions requiring fewer bits and flatter distributions requiring more (Fig. 1). These entropies rely solely on frequencies and therefore ignore a rich additional source of information about the system: the pairwise similarities and differences among its elements (Fig. 1).
@@
Figure 1: The concepts of element, frequency distribution, traditional entropy (here calculated at $q=1$), similarity function, similarity measure ($z_{ij}$, with examples), similarity matrix ($Z$), and sentropy (here calculated as LCR, also at $q=1$ for the most direct comparison). Entropy values are expressed in effective-number form, i.e. in units of the effective number of images present in the dataset. The choice of similarity measure is up to the investigator (see Section V.2); in this example, the similarity measure is defined as the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
@@
Incorporating similarity into the calculation results in so-called similarity-sensitive entropy or sentropy. This was first accomplished by Leinster and Cobbold [24] and extended by Reeve and colleagues [42], in a framework we refer to as LCR [27]. By incorporating similarity, LCR can differentiate between systems that have identical frequency distributions but whose elements vary in how alike they are (Fig. 1). This capability is valuable in many domains [37], especially in machine learning (ML) [11, 12] (for example, to achieve state-of-the-art performance more efficiently [10]), where datasets generally consist of all-unique elements, meaning any two datasets that are the same size will have the same traditional (i.e. similarity-insensitive) entropies because their frequency distributions are both flat, even though they may differ dramatically in composition. Sentropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [4, 5], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat, such that traditional entropies are less informative.
@@
More recently, an alternative form of sentropy to LCR has been described, consisting of the Vendi score [14, 40] and its variants or “cousins” [39]—hereafter collectively VS. This alternative raises the practical question of which form of sentropy to choose for a given system or dataset: LCR or VS. The present work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets commonly used for ML research, and comparing their values.
II Related Work
@@
Entropies fall into two classes: (i) traditional, i.e. similarity-insensitive, and (ii) similarity-sensitive, a.k.a. sentropy. Although the literature often refers to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), both classes are actually families of measures; individual members are distinguished by how they weight frequencies and, for sentropy, by their similarity matrix as well (see Table 1).
II.1 Traditional, similarity-insensitive entropy
@@
Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the dataset or system [18]. Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ down-weight less-frequent elements. $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form. Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for different $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [18]. For this reason, Hill’s formulation is often described as a unifying framework. $D_{q}$ are known as the Hill numbers or D numbers (for “diversity”) and are given by
$D_{q}(\mathbf{p})=\exp(H_{\alpha}(\mathbf{p}))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}^{q}\right)^{\frac{1}{1-q}},&q\neq 1\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln p_{i}\right),&q=1\end{cases}$ (1)
where $\mathbf{p}$ denotes the frequency distribution of unique elements. Other generalizations such as Tsallis entropy [47] also exist. What these traditional entropies have in common is that they are similarity-insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
II.2 Similarity-sensitive entropy: sentropy
II.2.1 The Leinster-Cobbold-Reeve framework (LCR)
LCR [27] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropies do not. A similarity matrix $Z$ is introduced, with entries $Z_{ij}\in[0,1]$ that quantify the similarity between elements $i$ and $j$. The similarity of each element to itself is set to 1, making $Z$’s diagonal entries 1. The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity-sensitive versions of the Rényi entropies $H_{\alpha}^{Z}$ and are given by
$\begin{split}D_{q}^{Z}(\mathbf{p};Z)&=\exp(H^{Z}_{\alpha}(\mathbf{p};Z))\\ &=\begin{cases}\left(\sum_{i=1}^{n}p_{i}(Z\mathbf{p})_{i}^{\,q-1}\right)^{\!\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\!\left(-\sum_{i=1}^{n}p_{i}\ln(Z\mathbf{p})_{i}\right),&q=1.\end{cases}\end{split}$ (2)
where $(Z\mathbf{p})_{i}=\sum_{j}z_{ij}p_{j}$ is the frequency-weighted average similarity of element $i$ to all the elements (including itself). When this average is large, element $i$ is considered ordinary, making Eq. 2 interpretable as the average “ordinariness” [24] across all elements. $D_{q}^{Z}{(\mathbf{p})}$ appears widely in the recent literature, where it is variously known as:
@@ II.2.1 The Leinster-Cobbold-Reeve framework (LCR)
  • similarity-sensitive or similarity-aware diversity more generally [37], for which “sentropy” is a convenient shorthand.
LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—i.e., where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become less informative (Fig. 1). Representative applications in the life sciences include high-throughput immunology (immunomes) [5], microbiome research (metagenomics) [37], and medical imaging [11]. Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (images), LCR has been shown to help identify performance predictors beyond dataset size or class balance [11].
II.2.2 The Vendi score and its variants (VS)
VS constitutes a related but separate class of sentropic measures. Like LCR, VS entropies are functions of a similarity matrix, but now the matrix has dimensions $n\times n$, where $n$ is the number of observations or total elements; for this reason we refer to it as $Z_{n}$ to distinguish it from the $Z$ used in LCR, which has one row/column per unique element. (When all elements are unique, $Z=Z_{n}$.) We can define $Z_{n}$ in terms of $Z$ as
@@ II.2.2 The Vendi score and its variants (VS)
The original VS is defined as the exponential of the Shannon entropy of the eigenvalues $\tilde{\lambda}_{i}=\frac{\lambda_{i}}{n}$ of $Z_{n}/n$ (where $\lambda_{i}$ are the eigenvalues of $Z$):
$\mathrm{VS}{(\mathbf{p}_{n};Z_{n})}=\exp\!\left(-\sum_{i=1}^{n}\tilde{\lambda}_{i}\log\tilde{\lambda}_{i}\right)$ (4)
The division by $n$ normalizes $Z_{n}$ to unit trace (because as in $Z$, the diagonals of $Z_{n}$ equal 1). VS is to this similarity matrix what the von Neumann entropy is to the quantum density matrix. Replacing the Shannon term in Eq. 4 with a Rényi entropy of order $q$ yields the so-called “cousins” [39] of the Vendi score:
$\mathrm{VS}_{q}{(\mathbf{p}_{n};Z_{n})}=\left(\sum_{i=1}^{n}\tilde{\lambda}_{i}^{q}\right)^{\frac{1}{1-q}}$ (5)
@@
$H_{\alpha}(\mathbf{p})=\frac{1}{1-\alpha}\log\sum_{i}p_{i}^{\alpha}$
@@
$q=1$ Yes $\tilde{\lambda}_{i}$ are eigenvalues of $Z_{n}=Z/N$ (positive semi-definite). ML diversity, Generative model evaluation, Ecology
@@
$\text{VS}_{q}(\mathbf{x},\mathbf{k})=\left(\sum_{i}\tilde{\lambda}_{i}^{q}\right)^{\tfrac{1}{1-q}}$
@@
III Theorem: VS bounds LCR
We now present our main theoretical result:
Theorem 1.
Let $Z$ be an $n\times n$ ($n>0$) positive semi-definite (PSD) similarity matrix, i.e. a PSD symmetric matrix with entries between 0 and $1$ whose diagonal entries equal $1$. Denote LCR and VS at order $q$ as $D_{q}(Z,p=\frac{1}{n})$ and $\text{VS}_{q}(Z)$, respectively, on the uniform distribution. Then for $q\geq 0$ (and for $q<0$ if $Z$ is full rank):
$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (7)
Following recent work [1, 17, 21, 49, 7], large language models (LLMs; especially Claude Opus 4.6) contributed to the proofs below. All results were manually checked.
Lemma 1.
Let $Z$ be an $n$-by-$n$ ($n>0$) non-negative symmetric matrix. Denote by $\lambda_{i}$ the (ordered) eigenvalues of $Z$, define $r_{i}=\sum_{j}Z_{ij}$, and let $R$ be the diagonal matrix defined by $R_{ii}=r_{i}$. Then $Z\preceq R$, i.e. RZR-Z is PSD.
Proof.
We must show that for any complex valued vector $v$, $v^{\dagger}\left(R-Z\right)v\geq 0$. Computing, we find
$v^{\dagger}\left(R-Z\right)v=\sum_{i}r_{i}|v_{i}|^{2}-\sum_{i,j}v_{i}^{\dagger}Z_{ij}v_{j}$ (8)
$=\sum_{i,j}Z_{ij}\left(|v_{i}|^{2}-v_{i}^{\dagger}v_{j}\right)$ (9)
$=\frac{1}{2}\sum_{i,j}\left[Z_{ij}\left(|v_{i}|^{2}-v_{i}^{\dagger}v_{j}\right)+Z_{ji}\left(|v_{j}|^{2}-v_{j}^{\dagger}v_{i}\right)\right]$ (10)
$=\frac{1}{2}\sum_{i,j}Z_{ij}|v_{i}-v_{j}|^{2}\geq 0$ (11)
where we have used the symmetry of Z, and the final inequality follows since Z is non-negative, i.e. $Z_{ij}\geq 0$. ∎
Lemma 2.
Let Z be as in 1 and define $\lambda_{i}$, $r_{i}$, and $R$ w.r.t. $Z$ as in 1. Construct
$M=R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}.$ (12)
(Note that since each $r_{i}\geq 1$, $R$ is invertible). Then $M$ is a PSD matrix with the same rank as $Z$, all of whose eigenvalues fall in the range $[0,1]$.
Proof.
[…]
Note that $Z$, $R$, and $R^{-1/2}$ are real symmetric matrices, and hence Hermitian, and furthermore by assumption $Z$ is PSD. Consequently $M=R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}$ is also PSD, as follows. Consider arbitrary complex vector $v$. Because $\left(R^{-\frac{1}{2}}\right)^{\dagger}=R^{-\frac{1}{2}}$, it is true that
$v^{\dagger}Mv=v^{\dagger}R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}v=(R^{-\frac{1}{2}}v)^{\dagger}Z(R^{-\frac{1}{2}}v)$ (13)
But since $Z$ is PSD, we have
$v^{\dagger}Mv=(R^{-\frac{1}{2}}v)^{\dagger}Z(R^{-\frac{1}{2}}v)\geq 0.$ (14)
Since $v$ was arbitrary, this proves $M$ is PSD.
To show that the eigenvalues of $M$ are bounded above by 1, it suffices to show that $M\preceq I$, i.e. that IMI-M is PSD. This follows because $Z\preceq R$, as proved in 1, but $I-M=R^{-\frac{1}{2}}\left(R-Z\right)R^{-\frac{1}{2}}$. We thus find that all the eigenvalues of $M$ must be in $[0,1]$.
An alternative demonstration that the eigenvalues are bounded above by $1$ follows from the Perron-Frobenius theorem [41, 15]. Since $M$ is non-negative, the Perron-Frobenius theorem guarantees that $M$ has a (not necessarily unique) non-negative eigenvector, and that all non-negative eigenvectors have a single eigenvalue corresponding to the spectral radius of $M$. Such an non-negative eigenvector is provided by $R^{\frac{1}{2}}\mathbf{1}$ (with $\mathbf{1}$ being the all-ones vector), which we confirm by computing
$MR^{\frac{1}{2}}\mathbf{1}=R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}R^{\frac{1}{2}}\mathbf{1}=R^{-\frac{1}{2}}Z\mathbf{1}=R^{-\frac{1}{2}}r=R^{\frac{1}{2}}\mathbf{1}$ (15)
Since the associated eigenvalue is $1$, and since we have already shown that $M$ is PSD, we find that all the eigenvalues of $M$ must be in $[0,1]$. ∎
Lemma 3.
Let $Z$ be as in 1. Then for real $q\geq 1$,
[…]
Likewise, for real $0<q<1$,
[…]
Proof.
Let $M$ be as in 2. 2 established that $M$ is PSD and has eigenvalues in $[0,1]$. Note that $Z=R^{\frac{1}{2}}MR^{\frac{1}{2}}$. By the Araki-Lieb-Thirring inequality [28, 3, 6], for any PSD matrices $A$ and $B$, and for any $p\geq 0$ and $q\geq 1$, we have:
[…]
Substituting $A=R^{\frac{1}{2}}$, $B=M$, and $p=1$, for $q>1$:
[…]
where the equality is by the cyclicity of the trace. However, because the eigenvalues of $M$ are in the range $[0,1]$ we have for $q>1$ that $M^{q}\preceq M$, meaning
[…]
proving the claim for $q>1$. For $0<q<1$, the direction of the Araki-Lieb-Thirring inequality reverses, yielding
[…]
but also $M\preceq M^{q}$ (since $x^{q}\geq x$ for $0\leq x\leq 1$ and $q\leq 1$) yielding
[…]
proving the claim for $0<q<1$. ∎
Proof of 1.
We can now prove 1 case by case.
Case 1: $q=+\infty$
The proof is an application of the Gershgorin circle inequality. First note that
$\mathrm{VS}_{\infty}(Z)$ $=\frac{n}{\lambda_{\text{max}}}$ (23)
$D_{\infty}(Z,\frac{1}{n})$ $=\frac{n}{r_{\text{max}}}$ (24)
where $\lambda_{\text{max}}$ is the largest eigenvalue of $Z$, i.e. $\lambda_{\text{max}}=\displaystyle\max_{i}(\lambda_{i})$, and likewise $r_{\text{max}}=\displaystyle\max_{i}(r_{i})$. However, by the Gershgorin circle inequality, every eigenvalue of $Z$ must fall within a Gershgorin interval, i.e. in one of the intervals $Z_{ii}-\sum_{j\neq i}Z_{i,j},Z_{ii}+\sum_{j\neq i}Z_{i,j}]$. As such, noting that $Z_{ii}+\sum_{j\neq i}Z_{i,j}=\sum_{j}Z_{i,j}=r_{i}$, we must have that every eigenvalue $\lambda_{\mu}$ of $Z$ is bounded by $\lambda_{\mu}\leq r_{\text{max}}$. But from 3, this means that
Case 2: $q>1$ and $0<q<1$
We now compute
$D_{q}(Z,\frac{1}{n})$ $=\left(\sum_{i=1}^{n}\frac{1}{n}(\sum_{j}Z_{ij}\frac{1}{n})^{q-1}\right)^{\frac{1}{1-q}}$
[…]
But since $f(x)=x^{\frac{1}{1-q}}$ is monotonically decreasing for $q>1$ (and $\frac{1}{n^{q}}$ is a positive constant), by 3 we get
[…]
On the other hand, since $f(x)=x^{\frac{1}{1-q}}$ is monotonically increasing for $0<q<1$, but the inequality from 3 reverses in this case, we get the same inequality in this case.
Case 3: $q=1$
We have proven the inequality for $q>1$ and $0<q<1$. Because both sides of the inequality are continuous functions of $q$ [27] [39], the result for $q=1$ follows by continuity.
Case 4: $q=0$
We begin by computing
$D_{0}(Z,\frac{1}{n})=\sum_{i}\frac{1}{r_{i}}$ (28)
[…]
Construct $M$ as in 2. The entries of $M$ are given by
$M_{ij}=\frac{Z_{ij}}{\sqrt{r_{i}r_{j}}},$ (30)
with diagonal entries
$M_{ii}=\frac{Z_{ii}}{r_{i}}=\frac{1}{r_{i}}.$ (31)
Consequently,
[…]
proving 1 for $q=0$, completing the proof for all $q\geq 0$.
Case 5: $q\in[-\infty,0)$ for $Z$ full rank
[…]
$VS_{q}(Z)\geq VS_{0}(Z)=n$ (34)
Finally, for $q=$-$\infty$, [27] proves that the LCR diversities likewise decrease monotonically in $q$, implying
$D_{-\infty}(Z,\frac{1}{n})\geq D_{q}(Z,\frac{1}{n})$ (35)
$D_{-\infty}(Z,\frac{1}{n})=\frac{n}{\displaystyle\min_{i}r_{i}}\leq n$ (36)
thereby completing the proof.
IV Experimental tests
IV.1 Datasets and definitions